OpenStudy (anonymous):

Parametric Differentiation x= cos^3 theta, y = sin^3 theta theta = pi/4 (45 degrees) Find dy/dx and evaluate the specified value of the parameter

OpenStudy (anonymous):

ok i just want to check my answer i got \[\sqrt{3}/12\]

OpenStudy (anonymous):

Let me get a piece of paper, 2 seconds

OpenStudy (anonymous):

\[dy/dx = 3\cos \theta (\sin \theta)^{2}/-3\sin (\cos \theta )^{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sqrt{3}/-6\]

OpenStudy (anonymous):

\[dy/dx=(3\sin^2(\theta)\cos(\theta))/(-3(\cos^2(\theta)\sin(\theta))\]

OpenStudy (anonymous):

evaluating I get -1

OpenStudy (anonymous):

u have your derivatives mixed the cos is squared in the numerator

OpenStudy (anonymous):

or do it? 1 sec

OpenStudy (anonymous):

\[dy/dx=-\sin(\theta)/\cos(\theta)=-\tan(\theta)\rightarrow-\tan(\pi/4)=-1\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i see what i did wrong let me do it over

OpenStudy (anonymous):

x = fn (theta) and y = fn (theta) hence dy/dx = [dy/d(theta)]/[dx/d(theta)] [dy/d(theta)] = 3sin^2 theta * cos theta [dx/d(theta)] = -3 cos^2 theta * sin theta dy/dx = -sin theta/cos theta = - tan theta since theta = pi/4 and tan pi/4 = 1 dy/dx = -1

OpenStudy (anonymous):

better explanation^^

OpenStudy (anonymous):

tell me if that doesn't make sense and I'll try to explain it a little better

OpenStudy (anonymous):

how'd you guys get tan theta?

OpenStudy (anonymous):

cancel out the cosine and the sine

OpenStudy (anonymous):

oki i got it thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

yay im the first person to give u a good answer! which it was indeed